OpenStudy (hybrik):

Find g(x), g(x^2) = 1/(x+3)

OpenStudy (hybrik):

\[g(x^2)=1/(x+3)\]

OpenStudy (hybrik):

g(x)=?

OpenStudy (hybrik):

@ganeshie8

OpenStudy (owlcoffee):

A very useful tip for this would be the equivalence exponents have: \[g(x)=g(\sqrt{x^2})\]

OpenStudy (hybrik):

So your saying to take the square root of the equation

OpenStudy (hybrik):

@Owlcoffee

OpenStudy (hybrik):

Anyone?

OpenStudy (hybrik):

?

OpenStudy (anonymous):

hi

OpenStudy (hybrik):

hi

OpenStudy (hybrik):

Anyone can help?

OpenStudy (anonymous):

i can try

OpenStudy (hybrik):

ok